OpenStudy (anonymous):

Solve by elimination 9x - 3y = 24 7x - 3y = 20 I got (6,-1) is that correct

OpenStudy (saifoo.khan):

Nopsie.

OpenStudy (anonymous):

No soorry that isnt't correct, do you want to know the correct answer and how to do it?

OpenStudy (psymon):

Start by multiplying the bottomequation by -1 and see what you can get from there :P

OpenStudy (anonymous):

First multiply the second equation by -1 and add it to the first equation. By doing this the y value will be eliminated

OpenStudy (anonymous):

So: 9x-3y=24 -7x+3y=-20 = 2x+0=4 x=2

OpenStudy (anonymous):

lol thx

OpenStudy (anonymous):

Now to solve for y you first substitute the x value into either equation 9(2)-3y=24 18-3y=24 -3y=6 y=-2 @Jessica_Love get it?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

@AwesomeB can you give me a medal

OpenStudy (anonymous):

ty for the medal :D

OpenStudy (anonymous):

np

OpenStudy (anonymous):

thx for medal :)

OpenStudy (anonymous):

Lol np